please help me solve this equation: The sum of two numbers is 8.5, their product is 18. Find the numbers..thanks very much
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

please help me solve this equation: The sum of two numbers is 8.5, their product is 18. Find the numbers.
----------------------------------------
x+y = 8.5
xy = 18
------------------
Solve using substitution:
x = 8.5-y
---
Substitute into xy = 18:
(8.5-y)y = 18
-y^2 + 8.5y = 18
y^2 - 8.5y + 18 = 0
(y-4.5)(y-4) = 0
---
If y = 4.5, x = 4
If y = 4, x = 4.5
========================
Cheers,
Stan H.

Answer by mananth(16946)

let one number be x
the other number is 8.5-x
..
x(8.5-x)=18
8.5x-x^2=18
17x/2 -x^2=18
17x^2-2x^2=36
2x^2-17x+36=0
2x^2-8x-9x+36=0
2x(x-4)-9(x-4)=0
(2x-9)(x-4)=0
x=9/2 OR 4
The numbers ar 4 & 4.5